A beam of monochromatic light of wavelength 4.0*10-7m in air passes into glass of refractive index there for the speed of light in a air is 3*10m/s calculate 1 the speed of light in glass 2 the frequency of light 3 the wavelength of light in glass


1
Expert's answer
2021-03-16T11:20:47-0400

(a) By the definition of the refractive index, we have:


"n=\dfrac{c}{v},""v=\dfrac{c}{n}=\dfrac{3.0\cdot10^8\ \dfrac{m}{s}}{1.5}=2.0\cdot10^8\ \dfrac{m}{s}."

b)

"f=\dfrac{c}{\lambda}=\dfrac{3\cdot10^8\ \dfrac{m}{s}}{4\cdot10^{-7}\ m}=750\ THz."

c)

"\lambda=\dfrac{\lambda_0}{n}=\dfrac{4\cdot10^{-7}\ m}{1.5}=2.67\cdot10^{-7}\ m."